### v2

# Masses and radiative decay widths of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ and their bottom analogs

Abstract
We study the mass spectra and radiative decays of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ in an unquenched framework. In addition to coupled channel effects between the $c\bar{s}$ cores and $D^{(*)}K$ channels, $D^{(*)}K$-$D^{(*)}K$ self interactions are also considered in this work and we succeed to reproduce their mass spectra. Furthermore, we study the radiative decays of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ by simultaneously including the compound structures of conventional $c\bar{s}$ cores and $D^{(*)}K$ components. We also calculate their bottom analogs with heavy quark symmetry. Our study offers useful insights into the important unquenched effects in the formation of $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$ and the bottom counterparts.
I Introduction
The rich hadron spectroscopy provides a good opportunity to decode the behavior of QCD in the nonperturbative region Liu:2024uxn ; Chen:2016spr ; Chen:2022asf ; Dong:2021juy ; Meng:2022ozq . With the abundance of hadron data, many exotic states have been discovered. Their decay behavior will help us disclose the nonperturbative strong interactions furthermore.
In 2003, the BABAR Collaboration first observed the $D_{s0}^{*}(2317)$ through the $D_{s}^{+}\pi^{0}$ invariant mass spectrum and estimated that its mass is near $2.32$ GeV and its width is small BaBar:2003oey . Shortly afterward the CLEO Collaboration confirmed the existence of $D_{s0}^{*}(2317)$ and found another state $D_{s1}^{\prime}(2460)$ in the $D_{s}^{+}\pi^{0}$ channel CLEO:2003ggt . They were further confirmed later by Belle and BABAR Belle:2003kup ; Belle:2003guh ; BaBar:2004yux ; BaBar:2003cdx ; BaBar:2006eep . Their isospin and spin parity quantum number $I(J^{P})$ are determined as 0($0(0^{+})$ and $0(1^{+})$ ParticleDataGroup:2024cfk , and their masses and widths are ParticleDataGroup:2024cfk
$\begin{split}m_{D_{s0}^{*}}(2317)&=2317.7\pm 0.5\;{\rm MeV},\\ \Gamma_{D_{s0}^{*}}(2317)&<3.8\;{\rm MeV},\\ m_{D_{s1}^{\prime}}(2460)&=2459.5\pm 0.6\;{\rm MeV},\\ \Gamma_{D_{s1}^{\prime}}(2460)&<3.5\;{\rm MeV}.\end{split}$
However, such experimental values of the masses of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are clearly smaller than the predictions from the conventional quark model Godfrey:1985xj ; Godfrey:2015dva ; Godfrey:1986wj ; Ebert:2009ua , which is usually referred as the low mass problem. This low mass problem also occurs in the well-known exotic hadrons $\Lambda(1405)$, $X(3872)$, and $\Lambda_{c}(2940)$ Silvestre-Brac:1991qqx ; vanBeveren:2003kd ; Kalashnikova:2005ui ; Li:2009ad ; Danilkin:2010cc ; Luo:2019qkm . Given the ongoing debate and various theoretical interpretations, the nature of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ and their bottom analogs remains an intriguing topic for further investigation.
Indeed, the discovery of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ states has sparked numerous theoretical works attempting to interpret their nature. In the modified quark model, the mass spectra of $D_{s}$ mesons were studied by including the screening effects Gao:2022bsb ; Song:2015nia . To explain the low mass problem, the $cq\bar{s}\bar{q}$ tetraquarks with open charms were investigated by using the constituent quark model, QCD sum rule, and so on Cheng:2003kg ; Dmitrasinovic:2005gc ; Maiani:2004vq ; Wang:2006uba ; Chen:2004dy ; Kim:2005gt ; Nielsen:2005ia . These two charmed mesons can also be the mixture of conventional $P$-wave quark-antiquark and four-quark components Vijande:2006hj .
Given that their masses are close to the $D^{(*)}K$ threshold, they are considered to be $S$-wave isoscalar $D^{(*)}K$ molecules in many phenomenological studies Kolomeitsev:2003ac ; Szczepaniak:2003vy ; vanBeveren:2003kd ; Barnes:2003dj ; Hofmann:2003je ; Gamermann:2006nm ; Guo:2006fu ; Guo:2006rp ; Flynn:2007ki ; Faessler:2007gv ; Guo:2009ct ; Xie:2010zza ; Liu:2011mi ; Cleven:2010aw ; Wu:2011yb ; Guo:2015dha ; Albaladejo:2016hae ; Du:2017ttu ; Guo:2018tjx ; Albaladejo:2018mhb ; Wu:2019vsy ; Kong:2021ohg ; Wang:2012bu ; Huang:2021fdt ; Ikeno:2023ojl ; Liu:2023uly ; Wu:2022wgn ; Liu:2022dmm ; Chen:2016ypj ; Chen:2022svh ; Liu:2024xbw ; Kim:2023htt ; Montesinos:2024uhq . The lattice QCD simulations of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ have also revealed the molecular composition Mohler:2011ke ; Liu:2012zya ; Mohler:2013rwa ; Lang:2014yfa ; Bali:2017pdv ; Alexandrou:2019tmk . The coupled channel effects between the $c\bar{s}$ cores and the $D^{(*)}K$ channels are found to be important for these two charmed mesons Simonov:2004ar ; Luo:2021dvj ; Hao:2022vwt ; Ni:2021pce ; Yang:2023tvc ; Coito:2011qn ; Hwang:2004cd . In addition, the $D^{(*)}K$ interactions were also considered in Refs. Ortega:2016mms ; Yang:2021tvc . The $D_{s0}^{*}(2317)$ is found to be dominated by the $DK$ channel with the Hamiltonian effective field theory in Ref. Yang:2021tvc , while the primary component of $D_{s0}^{*}(2317)$ is advocated to be the $c\bar{s}$ core in the resonating group method Ortega:2016mms .
Since heavy quark flavor symmetry establishes a connection between the open charm and bottom sectors, there has been considerable interest in predicting the bottom analogs of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$, i.e., $B_{s0}^{*}(1P)$ and $B_{s1}^{\prime}(1P)$. However, their bottom analogs are still not observed in experiments. The $B_{s}$ spectrum were studied with the conventional quark model Godfrey:2016nwn ; DiPierro:2001dwf ; Lu:2016bbk ; Ebert:2009ua ; Sun:2014wea ; li:2021hss . In Ref. Hao:2022ibj , the screened nonrelativistic quark model was used to investigate the masses of the $B_{s}$ mesons. The $B_{s0}^{*}(1P)$ and $B_{s1}^{\prime}(1P)$ were proposed as the two-antiquark and four-quark mixtures in Ref. Vijande:2007ke . In addition, the $B^{(*)}\bar{K}$ hadron molecule was advocated in Refs. Kolomeitsev:2003ac ; Guo:2006fu ; Guo:2006rp ; Cleven:2010aw ; Colangelo:2012xi ; Altenbuchinger:2013vwa ; Sun:2018zqs ; Ni:2021pce ; Kim:2023htt . The unquenched picture was adopted in Refs. Albaladejo:2016ztm ; Yang:2022vdb ; Ortega:2016pgg . The lattice QCD simulations were also carried out for these two states Gregory:2010gm ; Lang:2015hza ; Hudspith:2023loy .
Apart from the mass spectra, the radiative decays are also crucial to revealing the internal structures of these states. The radiative decays of $X(3872)$ were measured by different experimental groups Belle:2011wdj ; BaBar:2008flx ; LHCb:2014jvf ; BESIII:2020nbj ; LHCb:2024tpv and studied within various theoretical contexts Badalian:2012jz ; Dong:2009uf ; Cincioglu:2016fkm ; Chen:2024xlw . The E1 and M1 processes of charmonia and bottomia have been investigated in Refs. Godfrey:2015dia ; Segovia:2016xqb ; Wang:2018rjg ; Li:2012vc ; Deng:2016ktl ; Deng:2016stx , which can be compared with the experimental values listed in PDG ParticleDataGroup:2024cfk . The radiative decays of $D^{*}(2010)\to D\gamma$ were also measured experimentally CLEO:1992xqa ; BESIII:2014rqs .
The radiative transitions for conventional $D_{s}$ or $B_{s}$ mesons were explored in Refs. Godfrey:2003kg ; Godfrey:2016nwn ; Lu:2016bbk ; Godfrey:2015dva ; Li:2019tbn ; Radford:2009bs ; Chen:2020jku ; Green:2016occ ; Godfrey:2005ww ; Colangelo:2005hv ; Lutz:2007sk , and those for the conventional quark-antiquark and four-quark mixtures were studied in Refs. Vijande:2006hj ; Vijande:2007ke . Within the molecular picture, the processes of $D_{s0}^{*}(2317)\to D_{s}^{*}\gamma$, $D_{s1}^{\prime}(2460)\to D_{s}^{*}\gamma$, $D_{s1}^{\prime}(2460)\to D_{s0}^{*}(2317)\gamma$ were studied in Refs. Cleven:2014oka ; Fu:2021wde ; Faessler:2007gv ; Faessler:2007us ; Xiao:2016hoa ; Feng:2012zzf ; Feng:2012zze and $B_{s0}^{*}(1P)\to B_{s}^{*}\gamma$, $B_{s1}^{\prime}(1P)\to B_{s}^{*}\gamma$, $B_{s1}^{\prime}(1P)\to B_{s}\gamma$ were addressed in Refs. Faessler:2008vc ; Wang:2019ehs ; Cleven:2014oka ; Fu:2021wde . In this work, we will push forward the calculation of the radiative decays of $D_{s}$ and $B_{s}$ excitations by simultaneously including the conventional two-quark cores and $D^{(*)}K$/$B^{(*)}K$ components. We consider not only the coupled-channel effects between the bare $c\bar{s}$ cores and $D^{(*)}K$ channels but also the $D^{(*)}K$-$D^{(*)}K$ interactions. By solving the coupled-channel Schrödinger equations, the masses and the wave functions can be obtained. With such wave functions, we further deal with their radiative decays.
This article is organized as follows. After the introduction, we present the general formalism of the coupled-channel Schrödinger equation for the masses of the compound states in Sec. II, and the involved interactions are given in detail in Sec. III. We analyze the unquenched effect for $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom partners in Sec. IV. Their radiative decays are then studied in Sec. V. A short summary follows in Sec. VI.
II The coupled-channel Schrödinger equation
In this section, we will present the coupled-channel Schrödinger equation, which includes not only the coupling between the bare state and the hadron-hadron channel but also the self interaction of the two-hadron system.
It is plausible that the $D_{s0}^{*}(2317)$ could contain a single $c\bar{s}$ core plus the $DK$ cloud. To uncover this state, we need to set up the coupled-channel Schrödinger equation by including one bare state, which can be found in our previous work Zhang:2022pxc .
The $D_{s1}^{\prime}(2460)$ is a $J^{P}=1^{+}$ state and thus is related to $1P$ $c\bar{s}$ cores in the traditional quark model. There are two independent such cores $|c\bar{s},1^{+},j_{\ell}=1/2\rangle$ and $|c\bar{s},1^{+},j_{\ell}=3/2\rangle$ where $j_{\ell}$ is the angular momentum for the light quark freedom. To understand the $D_{s1}^{\prime}(2460)$, we need to solve the coupled-channel Schrödinger equation including two bare states. We will elaborate this formalism in the following.
The physical particle involving two bare states can be expressed as Weinberg:1965zz ; Guo:2017jvc ; Tornqvist:1995kr ; Kalashnikova:2005ui ; Danilkin:2009hr ; Eichten:1978tg ; Lu:2017hma ; Anwar:2018yqm ; Ortega:2021fem ; Ortega:2021yis ; Ortega:2016pgg ; Ortega:2009hj
$\begin{split}\left|\Psi\right\rangle=&c_{\alpha_{1}}\left|\Psi_{\alpha_{1}}\right\rangle+c_{\alpha_{2}}\left|\Psi_{\alpha_{2}}\right\rangle+\int{\rm d}^{3}\textbf{p}~{}\phi_{BC}(\textbf{p})\left|BC,\textbf{p}\right\rangle,\end{split}$ (1)
with the normalization condition
$|c_{\alpha_{1}}|^{2}+|c_{\alpha_{2}}|^{2}+\int\left|\phi_{BC}(\textbf{p})\right|^{2}{\rm d}^{3}{\bf p}=1,$ (2)
where the $\alpha_{1}$, $\alpha_{2}$, and $BC$ denote the first bare state, second bare state, and hadron-hadron channel, respectively. The full coupled-channel Schrödinger equation which contains two bare states can be then written as
$\left(\begin{array}[]{cccc}\hat{H}_{0}&\hat{H}_{0}&\hat{H}_{I}\\ \hat{H}_{0}&\hat{H}_{0}&\hat{H}_{I}\\ \hat{H}_{I}&\hat{H}_{I}&\hat{H}_{BC}\end{array}\right)\left(\begin{array}[]{c}c_{\alpha_{1}}|\Psi_{\alpha_{1}}\rangle\\ c_{\alpha_{2}}|\Psi_{\alpha_{2}}\rangle\\ \phi_{BC}(\textbf{p})\left|BC,\textbf{p}\right\rangle\end{array}\right)=M\left(\begin{array}[]{c}c_{\alpha_{1}}|\Psi_{\alpha_{1}}\rangle\\ c_{\alpha_{2}}|\Psi_{\alpha_{2}}\rangle\\ \phi_{BC}(\textbf{p})\left|BC,\textbf{p}\right\rangle\end{array}\right),$ (3)
where $\hat{H}_{0}$ is the Hamiltonian of the bare $c\bar{s}$ cores and can be obtained from the conventional potential model, $\hat{H}_{I}$ is the Hamiltonian describing the coupled-channel effect between the bare states and the $D^{*}K$ channel. As a novelty, we also consider $\hat{H}_{BC}$ for the $D^{*}K$-$D^{*}K$ interaction.
From Eq. (3), one obtains
$\begin{split}&M_{1}c_{\alpha_{1}}+M^{\prime}c_{\alpha_{2}}+\int\phi_{BC}({\bf p})H_{\alpha_{1}\to BC}^{*}({\bf p}){\rm d}^{3}{\bf p}=Mc_{\alpha_{1}},\\ &M^{\prime}c_{\alpha_{1}}+M_{2}c_{\alpha_{2}}+\int\phi_{BC}({\bf p})H_{\alpha_{2}\to BC}^{*}({\bf p}){\rm d}^{3}{\bf p}=Mc_{\alpha_{2}},\\ &c_{\alpha_{1}}H_{\alpha_{1}\to BC}({\bf p})+c_{\alpha_{2}}H_{\alpha_{2}\to BC}({\bf p})\\ &\qquad+\int\langle BC,{\bf p}|\hat{H}_{BC}\phi_{BC}({\bf p}^{\prime})|BC,{\bf p^{\prime}}\rangle{\rm d}^{3}{\bf p^{\prime}}=M\phi_{BC}({\bf p})\,,\end{split}$ (4)
where
$\begin{split}&M_{1}=\langle\Psi_{\alpha_{1}}|\hat{H}_{0}|\Psi_{\alpha_{1}}\rangle,\\ &M_{2}=\langle\Psi_{\alpha_{2}}|\hat{H}_{0}|\Psi_{\alpha_{2}}\rangle,\\ &M^{\prime}=\langle\Psi_{\alpha_{1}}|\hat{H}_{0}|\Psi_{\alpha_{2}}\rangle=\langle\Psi_{\alpha_{2}}|\hat{H}_{0}|\Psi_{\alpha_{1}}\rangle,\\ &H_{\alpha_{1}\to BC}({\bf p})=\langle BC,{\bf p}|\hat{H}_{I}|{\Psi_{\alpha_{1}}}\rangle,\\ &H_{\alpha_{2}\to BC}({\bf p})=\langle BC,{\bf p}|\hat{H}_{I}|{\Psi_{\alpha_{2}}}\rangle.\end{split}$ (5)
The matrix elements sandwiching the $\hat{H}_{BC}$ can be explicitly expressed as
$\begin{split}&\int\langle BC,{\bf p}|\hat{H}_{BC}\phi_{BC}({\bf p}^{\prime})|BC,{\bf p^{\prime}}\rangle{\rm d}^{3}{\bf p^{\prime}}\\ =&E_{BC}({\bf p})\phi_{BC}({\bf p})+\int V_{BC\to BC}({\bf p},{\bf p}^{\prime})\phi_{BC}({\bf p}^{\prime}){\rm d}^{3}{\bf p}^{\prime},\end{split}$ (6)
where $E_{BC}({\bf p})=m_{B}+m_{C}+\frac{p^{2}}{2m_{B}}+\frac{p^{2}}{2m_{C}}$ is the free energy of the intermediate $BC$ channel and $V_{BC\to BC}({\bf p},{\bf p}^{\prime})$ stands for the hadron-hadron interaction in momentum space.
To simplify the notations of Eq. (4), we define
$\begin{split}d_{1}=&\frac{M-M_{2}}{M^{2}-MM_{1}-{M^{\prime}}^{2}-MM_{2}+M_{1}M_{2}},\\ d_{2}=&\frac{M^{\prime}}{M^{2}-MM_{1}-{M^{\prime}}^{2}-MM_{2}+M_{1}M_{2}},\\ d_{3}=&\frac{M-M_{1}}{M^{2}-MM_{1}-{M^{\prime}}^{2}-MM_{2}+M_{1}M_{2}},\end{split}$ (7)
which give
$\begin{split}c_{\alpha_{1}}=&\int\phi_{BC}({\bf p})\big{[}d_{1}H_{\alpha_{1}\to BC}^{*}({\bf p})+d_{2}H_{\alpha_{2}\to BC}^{*}({\bf p})\big{]}{\rm d}^{3}{\bf p},\\ c_{\alpha_{2}}=&\int\phi_{BC}({\bf p})[d_{2}H_{\alpha_{1}\to BC}^{*}({\bf p})+d_{3}H_{\alpha_{2}\to BC}^{*}({\bf p})]{\rm d}^{3}{\bf p}\ .\end{split}$ (8)
Then Eq. (4) can be rewritten as
$\begin{split}&E_{BC}({\bf p})\phi_{BC}({\bf p})+\int{\rm d}^{3}{\bf p^{\prime}}\phi_{BC}({\bf p^{\prime}})\Big{[}V_{BC\to BC}({\bf p},{\bf p^{\prime}})\\ &\qquad+H_{\alpha_{1}\to BC}({\bf p})\left(d_{1}H_{\alpha_{1}\to BC}^{*}({\bf p^{\prime}})+d_{2}H_{\alpha_{2}\to BC}^{*}({\bf p^{\prime}})\right)\\ &\qquad+H_{\alpha_{2}\to BC}({\bf p})\left(d_{2}H_{\alpha_{1}\to BC}^{*}({\bf p^{\prime}})+d_{3}H_{\alpha_{2}\to BC}^{*}({\bf p^{\prime}})\right)\Big{]}\\ =&M\phi_{BC}({\bf p}).\end{split}$ (9)
Up to this point, we have provided a simple and effective method for solving the unquenched coupled-channel equations that include both two bare states and direct hadron-hadron channel.
We point out that if the relevant term of one of the bare states is removed, Eq. (LABEL:eq:FullChannel) is automatically reduced to the coupled-channel equation with only one bare state in our previous work Zhang:2022pxc . Furthermore, if we remove the relevant terms of both two bare states, the equation becomes a Schrödinger equation for the pure molecular state. We can also easily adjust the above formula to account for the situation when the interaction $V_{BC\to BC}({\bf p},{\bf p}^{\prime})$ of the direct hadron-hadron channel is dropped.
For obtaining the solution of Eq. (LABEL:eq:FullChannel), we use the complete-basis expansion method, where a set of the complete bases can be chosen as the harmonic oscillator basis, the Gaussian basis, and so on. In our study, we use the Gaussian basis to expand the wave function of bound hadron-hadron channel $\phi_{BC}(\textbf{p})$ in Eq. (LABEL:eq:FullChannel) Hiyama:2003cu ; Hiyama:2012sma
$\phi_{BC}({\bf p})=\sum_{i=1}^{N_{max}}C_{il}\phi_{ilm}^{p}({\bf p}),$ (10)
where $C_{il}$ is the coefficient of the corresponding basis, and $\phi_{ilm}^{p}({\bf p})$ denotes the Gaussian basis. In the coordinate space, the latter is given by
$\phi_{nlm}^{r}(\nu_{n},\textbf{r})=N_{nl}r^{l}e^{-v_{n}r^{2}}Y_{lm}(\hat{\textbf{r}}),$ (11)
where $N_{nl}$ is the normalization constant. By taking the Fourier transformation, the Gaussian basis in momentum space can be written as
$\phi_{nlm}^{p}(\nu_{n},\textbf{p})=(-i)^{l}\phi_{nlm}^{r}\left(\frac{1}{4\nu_{n}},\textbf{p}\right).$ (12)
In Eqs. (11) and (12), $v_{n}$ corresponds to the Gaussian ranges, i.e.,
$v_{n}=1/r_{n}^{2},\quad r_{n}=r_{1}a^{n-1}\quad(n=1,2\dots N_{\rm max}).$ (13)
In practice, we vary $a$, $r_{1}$ and $N_{\rm max}$ to make the eigenvalues corresponding to the ground states as low as possible while keeping the solutions stable enough, which gives $a=1.6238$, $r_{1}=0.0051$ fm and $N_{\rm max}=20$.
After introducing the above Gaussian basis, all the relevant Hamiltonian matrix elements can be expressed in simple forms:
$\begin{split}T_{fi}=&\int\textrm{d}^{3}\textbf{p}^{\prime}\phi_{flm}^{p*}(\nu_{f},\textbf{p}^{\prime})E_{BC}(\textbf{p}^{\prime})\phi_{ilm}^{p}(\nu_{i},\textbf{p}^{\prime}),\\ {\cal M}_{fi}=&\int\textrm{d}^{3}\textbf{p}^{\prime}\textrm{d}^{3}\textbf{p}\Big{[}H_{\alpha_{1}\to BC}({\bf p}^{\prime})\Big{(}d_{1}H_{\alpha_{1}\to BC}^{*}({\bf p})\\ &+d_{2}H_{\alpha_{2}\to BC}^{*}({\bf p})\Big{)}+H_{\alpha_{2}\to BC}({\bf p}^{\prime})\\ &\times\Big{(}d_{2}H_{\alpha_{1}\to BC}^{*}({\bf p})+d_{3}H_{\alpha_{2}\to BC}^{*}({\bf p})\Big{)}\Big{]}\\ &\times\phi_{flm}^{p*}(\nu_{f},\textbf{p}^{\prime})\phi_{ilm}^{p}(\nu_{i},\textbf{p}),\\ V_{fi}=&\int\textrm{d}^{3}\textbf{p}^{\prime}\textrm{d}^{3}\textbf{p}\phi_{flm}^{p*}(\nu_{f},\textbf{p}^{\prime})V_{BC\to BC}(\textbf{p},\textbf{p}^{\prime})\phi_{ilm}^{p}(\nu_{i},\textbf{p}),\\ N_{fi}=&\int\textrm{d}^{3}\textbf{r}^{\prime}\phi_{flm}^{r*}(\nu_{f},\textbf{r}^{\prime})\phi_{ilm}^{r}(\nu_{i},\textbf{r}^{\prime}).\end{split}$ (14)
With the above matrix elements, Eq. (LABEL:eq:FullChannel) can be recast into the following general eigenvalue equation
$\sum_{i=1}^{N_{max}}C_{il}(T_{fi}+{\cal M}_{fi}+V_{fi})=M\sum_{i=1}^{N_{max}}C_{il}N_{fi}\ .$ (15)
Because both sides of Eq. (15) depend on $M$, we are in fact dealing with a special eigenvalue-equation problem. We scan the possible values of $M$ in the left-side matrix $T+{\cal M}+V$ within a reasonable range until the eigenvalue of Eq. (15) gets equal to the input $M$, and this mass is exactly the solution of Eq. (15). The corresponding eigenvector gives the coefficient $C_{il}$.
III Detailed interactions
In this section, we provide the quark-model descriptions of the bare $c\bar{s}$/$\bar{b}s$ cores in Sec. III.1, the effective Lagrangian calculations of the $D^{(*)}K$/$B^{(*)}\bar{K}$ self-interactions in Sec. III.2 and the couplings between the bare states and the $D^{(*)}K/B^{(*)}\bar{K}$ in Sec. III.3, respectively. To implement these detailed interactions into the coupled-channel Schrödinger equations, one can then study the unquenched effects of the $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom analogs.
III.1 Quark-model description of the bare cores
We employ a nonrelativistic potential model to calculate the mass spectrum of the bare charmed-strange and bottom-strange states. The Hamiltonian $\hat{H}_{0}$ in Eq. (3) can be represented as Luo:2021dvj
$\hat{H}_{0}=\sum\limits_{i=1}\left(m_{i}+\frac{p_{i}^{2}}{2m_{i}}\right)+\sum\limits_{i<j}V_{ij},$ (16)
where $m_{i}$ and $p_{i}$ are the mass and momentum of the $i$-th constituent quark, respectively. $V_{ij}$ is the quark-quark/quark-antiquark interaction, which consists of the one-gluon-exchange potentials and confining potentials
$V_{ij}=H_{ij}^{\rm conf}+H_{ij}^{\rm hyp}+H_{ij}^{\rm so(cm)}+H_{ij}^{\rm so(tp)}.$ (17)
Here the first term denotes the spin-independent Cornell potential
$H_{ij}^{\rm conf}=-\frac{4}{3}\frac{\alpha_{s}}{r_{ij}}+br_{ij}+C,$ (18)
where $\alpha_{s}$, $b$, and $C$ stand for the coupling constant of one-gluon exchange, the strength of linear confinement, and the mass-renormalized constant, respectively. The hyperfine spin-spin interaction reads
$H_{ij}^{\rm hyp}=\frac{4\alpha_{s}}{3m_{i}m_{j}}\left(\frac{8\pi}{3}{\bf s}_{i}\cdot{\bf s}_{j}\tilde{\delta}(r)+\frac{1}{r_{ij}^{3}}S({\bf r}_{ij},{\bf s}_{i},{\bf s}_{j})\right),$ (19)
where
$\tilde{\delta}(r)=\frac{\sigma^{3}}{\pi^{3/2}}{\rm e}^{-\sigma^{2}r^{2}}$ (20)
is a Gaussian smearing function including a smearing parameter $\sigma$, and
$S({\bf r}_{ij},{\bf s}_{i},{\bf s}_{j})=\frac{3{\bf s}_{i}\cdot{\bf r}_{ij}{\bf s}_{j}\cdot{\bf r}_{ij}}{r_{ij}^{2}}-{\bf s}_{i}\cdot{\bf s}_{j}$ (21)
is a tensor operator. The color-magnetic term and Thomas-precession piece of the spin-orbit interactions could be expressed as
$H_{ij}^{\rm so(cm)}=\frac{4\alpha_{s}}{3r_{ij}^{3}}\left(\frac{1}{m_{i}}+\frac{1}{m_{j}}\right)\left(\frac{{\bf s}_{i}}{m_{i}}+\frac{{\bf s}_{j}}{m_{j}}\right)\cdot{\bf L}\,,$ (22)
and
$H_{ij}^{\rm so(tp)}=-\frac{1}{2r_{ij}}\frac{\partial H_{ij}^{\rm conf}}{\partial r_{ij}}\left(\frac{{\bf s}_{i}}{m_{i}^{2}}+\frac{{\bf s}_{j}}{m_{j}^{2}}\right)\cdot{\bf L}\,,$ (23)
respectively.
As the $D_{s}$ and $B_{s}$ are typical heavy-light systems, it is convenient to construct the wave function in the heavy quark symmetry, i.e.,
$\Psi_{JM}\sim[[s_{q}\phi_{NL}({\bf r})]_{j_{\ell}}s_{Q}]_{JM},$ (24)
where $s_{q}$ and $s_{Q}$ are the spins of the light and heavy quarks, respectively. $\phi_{NL}({\bf r})$ is the spatial wave function, and $N$ and $L$ are radial and orbital quantum numbers, in order.
By reproducing the masses of the low-lying well-established $\pi$, $K$, $D$, $D_{s}$, and $B_{s}$ mesons, we can determine the parameters in the quenched potential model, and the explicit results are: $m_{u/d}=0.336$ GeV, $m_{s}=0.540$ GeV, and $m_{c}=1.660$ GeV, $m_{b}=4.730$ GeV, $\alpha_{s}=0.570$, $b=0.144$ GeV2, $\sigma=1.028$ GeV, and $C_{D_{s}}=-0.447$ GeV, $C_{B_{s}}=-0.122$ GeV.
Using these parameters and the interactions above, we obtain the relevant bare masses in Table 1, which are close to the results in Ref. Ebert:2009ua .
Table 1: The quenched masses in the nonrelativistic potential model. Here, $M_{0}$ is the bare mass of $D_{s0}^{*}(2317)$. $M_{1}$ and $M_{2}$ are the bare masses of the $j_{\ell}=1/2$ and $3/2$ components in the $D_{s1}^{\prime}(2460)$, respectively, and $M^{\prime}$ is their transition matrix element as defined in Eq. (LABEL:eq:TranM). The masses are given in units of MeV.
Core $M_{0}$ $M_{1}$ $M_{2}$ $M^{\prime}$
$c\bar{s}$ 2447 2537 2526 4
$s\bar{b}$ 5821 5862 5861 2
III.2 Two-hadron interactions in effective Lagrangian
We employ the effective Lagrangian method to calculate the hadron-hadron interactions $V_{BC\to BC}$. Here we simultaneously consider the contact term and the exchange of light-mesons ($\sigma$, $\rho$, $\omega$) as illustrated in Fig. 1 to describe the interaction of the $S$-wave $D^{(*)}K$/$B^{(*)}\bar{K}$ system.
Figure 1: The Feynman diagrams for the $D^{(*)}K$/$B^{(*)}\bar{K}$ interaction. In the heavy quark limit, the interactions of $D^{*}K$ and $B^{(*)}\bar{K}$ share the same forms as $DK$.
[…]
The Lagrangian with the $DK$ contact term takes the form Wise:1992hn ; Manohar:2000dt
$\mathcal{L}_{\mathcal{D}\mathcal{D}\mathcal{P}\mathcal{P}}=\frac{i}{2f_{K}^{2}}\bar{\mathcal{D}}(\mathcal{P}v\cdot\partial\mathcal{P}-v\cdot\partial\mathcal{P}\mathcal{P})\mathcal{D}.$ (25)
Here, $v=(1,0,0,0)$ stands for the 4-velocity of $D$ meson, and $\mathcal{D}$ denotes the charmed meson fields $\mathcal{D}^{T}=(D^{0},~{}D^{+},D_{s}^{+})$. $f_{K}$ is the kaon decay constant with the value of $f_{K}=0.113$ MeV. The pseudoscalar field $\mathcal{P}$ is
$\mathcal{P}=\left(\begin{array}[]{ccc}\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&\pi^{+}&K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\frac{\eta}{\sqrt{6}}&K^{0}\\ K^{-}&\bar{K}^{0}&-\frac{2}{\sqrt{6}}\eta\end{array}\right)\,.$ (28)
The Lagrangians with the $DD\sigma$ and $DDV$ couplings are given by Yan:1992gz ; Burdman:1992gh ; Casalbuoni:1996pg ; Falk:1992cx
$\mathcal{L}_{{\mathcal{D}}{\mathcal{D}}\sigma}=-2g_{s}{\mathcal{D}}_{b}^{{\dagger}}{\mathcal{D}}_{b}\sigma\,,$ (29)
$\mathcal{L}_{{\mathcal{D}}{\mathcal{D}}{\mathcal{V}}}=-\sqrt{2}\beta g_{V}{\mathcal{D}}_{b}{\mathcal{D}}_{a}^{{\dagger}}v\cdot{\mathcal{V}}_{ba}\,,$ (30)
where the vector meson matrix $\mathcal{V}$ is
$\mathcal{V}=\left(\begin{array}[]{ccc}\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&\rho^{+}&K^{*+}\\ \rho^{-}&-\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}&K^{*0}\\ K^{*-}&\bar{K}^{*0}&\phi\end{array}\right).$ (33)
The above equations give a vanishing $DD\phi$ coupling since it is Okubo-Zweig-Iizuka (OZI) suppressed. The Lagrangian with the couplings of $K\bar{K}\sigma$, $K\bar{K}\omega$ and $KK\rho$ can be written as Lin:1999ad ; Nagahiro:2008mn
$\mathcal{L}_{\sigma KK}$ $=$ $g_{\sigma}m_{K}\bar{K}K\sigma,$ (34)
$\mathcal{L}_{\rho KK}$ $=$ $ig_{\rho KK}\left(\bar{K}\partial^{\mu}K-\partial^{\mu}\bar{K}K\right)\bm{\tau}\cdot\bm{\rho_{\mu}},$ (35)
$\mathcal{L}_{\omega KK}$ $=$ $ig_{\omega KK}\left(\bar{K}\partial^{\mu}K-\partial^{\mu}\bar{K}K\right)\omega_{\mu},$ (36)
where $g_{s}=g_{\sigma}$ is estimated by the quark model Wang:2019aoc . The $\sigma$ meson coupling constant $g_{s}$ can be calculated by the quark model: $g_{s}=g_{\sigma NN}/3$ with $g_{\sigma NN}^{2}/(4\pi)=5.69$ Isola:2003fh . Following Ref. Chen:2011cj , we take $g_{\rho{K}^{(*)}{K}^{(*)}}=3.425$ and $g_{\omega{K}^{(*)}{K}^{(*)}}=4.396$. Using the vector meson dominance, $\beta$ can be fixed as 0.9, and $g_{V}=m_{\rho}/f_{\pi}=5.8$ Sun:2011uh .
With these Lagrangians, we can calculate the Feynman amplitudes $\mathcal{M}_{DK\to DK}({\bf q})$ corresponding to the Feynman diagrams in Fig. 1. Then the effective potential related to such amplitudes can be given by the Breit approximation Breit:1929zz ,
$\mathcal{V}_{DK\to DK}({\bf q})=-\frac{\mathcal{M}_{DK\to DK}({\bf q})}{\sqrt{\Pi_{i}2m_{i}\Pi_{f}2m_{f}}},$ (37)
where $m_{i}$ and $m_{f}$ are the masses of the initial and final states, respectively. The flavor wave function of the isoscalar $D^{(*)}K$ system with $\left|I=0,I_{3}=0\right\rangle$ is
$\frac{1}{\sqrt{2}}\left(\left|D^{(*)+}K^{0}\right\rangle+\left|D^{(*)0}K^{+}\right\rangle\right).$ (38)
In momentum space, these effective potentials of the Feynman diagrams in Fig. 1 are found to be
$\begin{split}&V_{\rm{Contact}}=-\frac{m_{K}}{2f_{K}^{2}},\\ &V_{\rho}=-\frac{3}{2}\beta g_{V}g_{\rho KK}\frac{1}{{\textbf{q}}^{2}+m_{\rho}},\\ &V_{\omega}=-\frac{1}{2}\beta g_{V}g_{\omega KK}\frac{1}{{\bf q}^{2}+m_{\omega}^{2}},\\ &V_{\sigma}=-g_{s}g_{\sigma}\frac{1}{{\bf q}^{2}+m_{\sigma}^{2}},\\ &\mathcal{V}_{\rm Total}=V_{\rm{Contact}}+V_{\rho}+V_{\omega}+V_{\sigma}\,,\end{split}$ (39)
where the approximation of $q^{0}=0$ is taken.
Figure 2: The $I=0$ interaction potentials of $D^{(*)}K$/$B^{(*)}\bar{K}$ systems in coordinate space. Notice that with the heavy quark symmetry, the $D^{(*)}K$ and $B^{(*)}\bar{K}$ potentials are equal.
Based on the effective potential $\mathcal{V}_{BC\to BC}({\bf q})$ in momentum space, its counterpart $\mathcal{V}_{BC\to BC}({\bf r})$ in coordinate space can be obtained via the following Fourier transformation
$V_{BC\to BC}({\bf r})=\int\frac{\textrm{d}^{3}{\bf q}}{(2\pi)^{3}}{\rm e}^{i{\bf q}\cdot\textbf{r}}\mathcal{V}_{BC\to BC}({\bf q})\mathcal{F}({\bf q})\,,$ (40)
where $\mathcal{F}(\textbf{q})={\rm e}^{-\textbf{q}^{2n}/\Lambda^{2n}}$ is the form factor with Gauss form to get rid off divergences showing up in $|\textbf{q}|\to\infty$ limit and to also renormalize the potential Xu:2017tsr . The matrix element $V_{fi}$ in Eq. (14) can be then conveniently calculated via
$V_{fi}=\int\textrm{d}^{3}\textbf{r}\phi_{flm}^{r*}(\nu_{f},\textbf{r})V_{BC\to BC}(\textbf{r})\phi_{ilm}^{r}(\nu_{i},\textbf{r})\,.$ (41)
So far we have focused on the $DK$ effective potential. In the heavy-quark limit, the interactions of $D^{*}K$ and $DK$ system are in fact exactly the same. In Fig. 2, we illustrate their potentials in coordinate space. It is evident that the contribution from each term, including the contact interaction and the exchanges of $\rho$, $\sigma$ and $\omega$, is attractive. The effects of the $\sigma$ and $\omega$ exchanges are rather weak, with similar strengths. The strongest attractive interaction is provided by the $\rho$ exchange and it is dominant over the others. Furthermore, the potential with the cutoff $\Lambda=0.7$ GeV is found to be more attractive than that with $\Lambda=0.4$ GeV.
In the heavy quark limit, the $S$-wave $B^{(*)}\bar{K}$ interaction shares the same form as that of the $D^{(*)}K$ system Chen:2016ypj ; Chen:2022svh . Nevertheless, the binding properties of the two systems are different because of their different reduced masses $\mu_{AB}=M_{A}M_{B}/(M_{A}+M_{B})$.
III.3 Interaction between the bare core and hadron-hadron channel
For the interaction between the bare states and $D^{(*)}K$/$B^{(*)}\bar{K}$ in Eq. (LABEL:eq:FullChannel), we employ the quark-pair-creation model, whose Hamiltonian is given by Micu:1968mk ; LeYaouanc:1972vsx
$\hat{H}_{I}=g\int\textrm{d}^{3}x\bar{\psi}(x)\psi(x).$ (42)
Here, $g=2m_{q}\gamma$, $m_{q}$ is the mass of the creation quark, and the dimensionless parameter $\gamma$ describes the strength of the quark and antiquark pair creation from the vacuum, which can be determined phenomenologically by the OZI allowed decay widths of charmonia.
In the nonrelativistic limit, $\hat{H}_{I}$ is equivalent to Chen:2016iyi
$\begin{split}\hat{H}_{I}=&-3\gamma\sum_{m}\left<1,m;1,-m|0,0\right>\int\textrm{d}^{3}\textbf{p}_{\mu}d^{3}\textbf{p}_{\nu}\delta\left(\textbf{p}_{\mu}+\textbf{p}_{\nu}\right)\\ &\times\mathcal{Y}_{1}^{m}\left(\frac{\textbf{p}_{\mu}-\textbf{p}_{\nu}}{2}\right)\omega^{\left(\mu,\nu\right)}\phi^{\left(\mu,\nu\right)}\chi_{-m}^{\left(\mu,\nu\right)}b_{\mu}^{\dagger}\left(\textbf{p}_{\mu}\right)d_{\nu}^{\dagger}\left(\textbf{p}_{\nu}\right)\,,\end{split}$ (43)
where $\omega$, $\phi$, $\chi$, and $\mathcal{Y}_{1}^{m}$ denote the color, flavor, spin, and orbital angular momentum functions of the quark pair, respectively. $b_{\mu}^{\dagger}$ and $d_{\nu}^{\dagger}$ are quark and antiquark creation operators, in order. In this work, we determine the dimensionless parameter $\gamma$ to be $5.7$ from the fit to the total decay width of $D_{s2}^{*}(2573)$.
Using the wave functions to describe the bare hadrons in previous Sec. III.1, we can obtain the transition amplitudes of bare cores to hadron-hadron channels in Eq. (LABEL:eq:TranM) with the quark-pair-creation model.
IV Spectra from coupled Schrödinger equations
Based on the formalism described in Sec. II and the detailed interaction in Sec. III, we are ready to study the dynamical couplings between the $S$-wave $D^{(*)}K$ channels and the bare $c\bar{s}$ cores, and explore the unquenched effects for the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ mesons. Within the same framework, we will also investigate their bottom analogs.
IV.1 Quenched results from the bare $c\bar{s}$ cores
In Sec. III, we enumerate the results under the quenched picture. We can see that the bare mass of $|c\bar{s},0^{+},j_{\ell}=1/2\rangle$ is obtained to be $2447$ MeV, about 80 MeV above the $DK$ threshold and $130$ MeV larger than the measured mass of $D_{s0}^{*}(2317)$. The bare mass of $|c\bar{s},1^{+},j_{\ell}=1/2\rangle$ is obtained to be $2537$ MeV, which is about $32$ MeV above the $D^{*}K$ threshold and $77$ MeV larger than the measured mass of $D_{s1}^{\prime}(2460)$. Thus we can clearly see that the near-threshold behavior and the prominent low-mass problem exist in these two states.
From Table 1, the mixing interaction $M^{\prime}$ between the $|c\bar{s},1^{+},j_{\ell}=1/2\rangle$ and $|c\bar{s},1^{+},j_{\ell}=3/2\rangle$ is weak. In the heavy quark limit, $j_{\ell}$ is a good quantum number and different states with different $j_{\ell}$ will not mix.
IV.2 Adding the couplings between $c\bar{s}$ core and $D^{(*)}K$
Because the masses of the physical states $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are clearly lower than the bare $c\bar{s}$ cores, they cannot be explained as the $c\bar{s}$ pairs in the simple quenched quark model. In this part, we explore the situation by including the coupling between $c\bar{s}$ and $D^{(*)}K$ but excluding the $D^{(*)}K$-$D^{(*)}K$ interaction, and give the results in the middle part of Table 2.
According to the values in this table, the resulting masses decrease due to the couplings between $c\bar{s}$ core and $D^{(*)}K$ channel, and they are now closer to the measured values. In the $D_{s0}^{*}$ case, the probabilities of the bare $c\bar{s}$ core and $DK$ are $42\%$ and $58\%$, respectively. In the $D_{s1}^{\prime}$ case, the proportions of $D^{*}K$ and the bare states with $j_{\ell}=1/2$ and $j_{\ell}=3/2$ are $37.4\%$, $62.1\%$, and $0.5\%$, in order. We see that the component of $|c\bar{s},1^{+},j_{\ell}=1/2\rangle$ is dominant for $D_{s1}^{\prime}$.
Although the masses are shifted down, one cannot still fully reproduce the experimental masses of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$. Therefore, additional effects need to be considered.
Table 2: The masses in the quenched quark model, with coupled-channel effects between the bare cores and the $D^{(*)}K/B^{(*)}\bar{K}$, and with $D^{(*)}K$-$D^{(*)}K$/$B^{(*)}\bar{K}$-$B^{(*)}\bar{K}$ interactions furthermore. Here, $r_{\rm RMS}$ refers to the root-mean-square radius of the $D^{(*)}K/B^{(*)}\bar{K}$ component, and $P({\rm bare})$ represents the probability of bare $c\bar{s}/\bar{b}s$ core in the eigenstate.
Cases Quenched Without hadron-hadron interaction (approximation) With hadron-hadron interaction (full result)
[…]
$D_{s0}^{*}(2317)$ 2447 2349 1.84 42.0 2321 1.33 34.4
$D_{s1}^{\prime}(2460)$ 2537 2478 1.48 62.6 2456 1.30 47.8
$B_{s0}^{*}(1P)$ 5821 5741 1.30 58.2 5714 1.17 44.5
$B_{s1}^{\prime}(1P)$ 5862 5787 1.26 61.7 5762 1.16 47.2
IV.3 Further adding the $D^{(*)}K$-$D^{(*)}K$ interaction
In addition to the effects already discussed in Sec. IV.2, we further take into account the $D^{(*)}K$-$D^{(*)}K$ interactions in Eq. (LABEL:eq:FullChannel) to solve the unquenched effect of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$. We list the results after including the $D^{(*)}K$ interaction with the cutoff $\Lambda=0.41$ GeV in the right part of Table 2. The results show that the $D^{(*)}K$-$D^{(*)}K$ interaction can indeed further lower the masses down. Now the resulting masses from our study are very close to the experimental ones of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$.
For easy comparisons, the spectra of the quenched quark model and those with the full coupled-channel effects as shown in Table 2 are visualized in the left part of Fig. 3. The masses in the quenched quark model are much larger than the experimental ones. With the continuous improvement of the unquenched effects, the theoretical masses gradually overlap with the experimental ones. Both the $c\bar{s}$-$D^{(*)}K$ couplings and $D^{(*)}K$-$D^{(*)}K$ interactions are important to understand the inner structures of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$.
Figure 3: Comparison of the masses in Table 2 and experimental measurements. See Table 2 for the meaning of notations. The results labeled as this work correspond to the full result in our study.
Using Eqs. (8) and (11), we can obtain the $D^{(*)}K$ radial wave function and bare $c\bar{s}$ core probability $|c_{\alpha_{i}}|^{2}$ which can further help us to reveal the role of $D^{(*)}K$ channels and bare $c\bar{s}$ cores in physical state $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$. From Table 2, we can see the $0^{+}$ state, i.e., $D_{s0}^{*}(2317)$, is a little fatter with larger radius and contains lesser bare $c\bar{s}$ core than the $1^{+}$ one, which is consistent with the conclusions from earlier studies MartinezTorres:2014kpc ; Albaladejo:2018mhb ; Tan:2021bvl . The $D^{(*)}K$ possibility in the physical states $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are about $65.6\%$ and $52.2\%$, respectively. The predominantly $D^{(*)}K$ structure of these two states has an indirect robust theoretical support in Refs. MartinezTorres:2014kpc ; Albaladejo:2016lbb ; Du:2017zvv .
IV.4 Closer inspection of binding energy and radial wave function
If $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are pure $D^{(*)}K$ molecules, their binding energies would be about 50 MeV. They would be very deeply bound states compared to the deuteron, which indicates the attractive $D^{(*)}K$ interactions must be extremely larger than the $pn$ attraction! However, the $D^{(*)}K$-$D^{(*)}K$ interactions in our framework are not extraordinarily strong. If neglecting the $c\bar{s}$ core, the pure $DK$ and $D^{*}K$ systems can only form bound states with binding energies around $2$ MeV and $1$ MeV, respectively. This means the interaction strength for $D^{(*)}K$ in this work is comparable to that for $pn$ in the deuteron.
Of course, we can artificially increase the cutoff $\Lambda$ in order to obtain very large $D^{(*)}K$ interactions, so that one can obtain the masses of pure $D^{(*)}K$ molecules close to the experimental values. When the cutoff $\Lambda$ reaches $0.73$ GeV, the masses of the pure $0^{+}$ $DK$ and $1^{+}$ $D^{*}K$ become $2318$ MeV and $2458$ MeV, respectively. In this case, the interaction strength for $D^{(*)}K$ is very different from that of the $pn$ system. However, we do not exactly know how different their interaction strengths should be in the real world, and thus we also show the results for the pure molecule assumption in the following discussion.
In addition to the spectra, the bare state will also bring a correction to the $D^{(*)}K$ radial wave functions. In Fig. 4, we give the $D^{(*)}K$ radial wave functions in two cases: the pure $D^{(*)}K$ molecule (the red dot line) and the physical state dressed with $D^{(*)}K$ coupled channel involving hadron-hadron interaction (the black solid line). The results indicate that the bare states can obviously affect $D^{(*)}K$ radial wave functions and bind the $D^{(*)}K$ tighter. The radial wave functions will be used later in the calculation of the radiative decay of these two states.
IV.5 The prediction for the bottom analogs
With the method of describing the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$, we use the same coupling constants and cutoff parameters to further give the predictions of the spectra of their bottom analogs. Here, we consider their bottom-strange analogs as the $\bar{b}s$ cores plus $S$-wave $B^{(*)}\bar{K}$ channels.
According to the values of Table 2, the two $B_{s}(1P)$ states have approximate root-mean-square radius. The probability of bare core in $B_{s0}^{*}(1P)$ is bigger than that in $D_{s0}^{*}(2317)$.
From Fig. 3, the masses of the two $B_{s}(1P)$ states clearly decrease after considering the $B^{(*)}\bar{K}$ interactions. When using the cutoff that reproduces the masses of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$, we can predict $M_{B_{s0}^{*}}=5714$ MeV and $M_{B_{s1}^{\prime}}=5762$ MeV.
Figure 4: The $D^{(*)}K$ radial wave functions $u(r)=r\phi_{D^{(*)}K}(r)$ for the pure $D^{(*)}K$ molecule (with red dashed lines), and the dressed state with $D^{(*)}K$ interaction (with black solid lines).
V Radiative decay
To shed more light on the properties of $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, we further explore their electromagnetic properties. We systematically study the electromagnetic properties of the bare cores, the pure $D^{(*)}K$/$B^{(*)}\bar{K}$ molecular states, and their physical mixtures.
V.1 The E1 transitions for the bare cores
The E1 transitions of the bare $D_{s}$/$B_{s}$ states play important roles in their electromagnetic properties. Under the assumption that they are simple $c\bar{s}$ or $\bar{b}s$ states, the partial widths associated with the E1 transitions can be obtained using the following expression Brodsky:1968ea ; Kwong:1988ae
$\Gamma(i\to f+\gamma)=\frac{4}{3}\langle e_{Q}\rangle^{2}\alpha\omega^{3}C_{fi}\delta_{SS^{\prime}}|\langle f|r|i\rangle|^{2},$ (44)
with
$C_{fi}$ $=max\left(L,L^{\prime}\right)\left(2J^{\prime}+1\right)\left\{\begin{array}[]{ccc}L^{\prime}&J^{\prime}&S\\ J&L&1\end{array}\right\}^{2},$ (47)
and
$e_{Q}=\left\{\begin{array}[]{ll}(m_{s}e_{c}-m_{c}e_{\bar{s}})/(m_{c}+m_{s})&\quad\mbox{for}\;D_{s}\;\mbox{meson}\\ (m_{b}e_{s}-m_{s}e_{\bar{b}})/(m_{b}+m_{s})&\quad\mbox{for}\;B_{s}\;\mbox{meson}\end{array}\right..$ (48)
[…]
If the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are assumed to be conventional $c\bar{s}$ mesons with only quark-antiquark cores, the widths for $D_{s0}^{*}\to D_{s}^{*}\gamma$, $D_{s1}^{\prime}\to D_{s}^{*}\gamma$, and $D_{s1}^{\prime}\to D_{s}\gamma$ are predicted to be 12.7, 14.3, and 6.0 keV, respectively. For the bottom analogs, the widths for $B_{s0}^{*}\to B_{s}^{*}\gamma$, $B_{s1}^{\prime}\to B_{s}^{*}\gamma$, and $B_{s1}^{\prime}\to B_{s}\gamma$ are 73.3, 77.2, and 44.9 keV, correspondingly. The bottom mesons have larger widths because of the larger values of $e_{Q}$ in Eq. (48).
V.2 The radiative decay for pure molecular state
If the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are pure $D^{(*)}K$ molecular states, their radiative decays can be totally different from those of conventional quark-antiquark mesons. We study the processes that the pure $D^{(*)}K$ molecules decay to $D_{s}^{(*)}\gamma$ in this subsection.
We denote the molecular state consisting of hadron $A$ and hadron $B$ as $[AB]$, and the radiative decay widths $\Gamma_{[AB]\to C\gamma}$ can be calculated as follows. The decay amplitude is Luo:2023hnp ; Zhang:2006ix ; Chen:2024xlw
[$\begin{split}\mathcal{M}^{JM}_{[AB]\to C\gamma}=&\int\frac{\rm d^{3}\textbf{p}^{\prime}}{(2\pi)^{3/2}}\,\frac{\sqrt{2m_{[AB]}}}{\sqrt{2m_{A}}\sqrt{2m_{B}}}\hat{\phi}^{JM}_{[AB]}({\bf p^{\prime}})\\ &\otimes\hat{\mathcal{M}}_{AB\to C\gamma}({\bf p},{\bf p}^{\prime}),\\ \end{split}$ (49)
$\hat{\phi}^{JM}_{[AB]}({\bf p}^{\prime})$ $=$ $\bigg{\{}\phi_{[AB]|^{3}L_{J}\rangle}(|\textbf{p}^{\prime}|)\,\,C^{S,m_{S}}_{1m_{1},1m_{2}}C^{J,M}_{Sm_{S},Lm_{L}}\,Y_{L,m_{L}}(\theta,\phi)$ (50)
$\qquad\bigg{|}\forall~{}L,S,m_{L},m_{1},m_{2}\bigg{\}}.$
where ^$\hat{\mathcal{M}}_{AB\to C\gamma}({\bf p},{\bf p}^{\prime})$ is the scattering amplitude of $AB\to C\gamma$ process, and $\hat{\phi}^{JM}_{[AB]}({\bf p}^{\prime})$ denotes the wave function of the molecular state $[AB]$ in momentum space which can be obtained via Eq. (15).
Then the decay width can be expressed as
$\Gamma_{[AB]\to C\gamma}=\frac{1}{2J+1}\frac{|\textbf{p}|}{32\pi^{2}m^{2}_{\left[AB\right]}}\sum_{M}\int\left|\mathcal{M}^{JM}_{[AB]\to C\gamma}\right|^{2}{\rm d}\Omega_{\textbf{p}}\,,$ (51)
where the factor $1/(2J+1)$ comes from the spin average over the initial polarization vector. In Eqs. (49)-(51), ${\bf p}^{\prime}$ and ${\bf p}$ are the momentum of $A$ and $C$, respectively.
To give the radiative decay width $\Gamma_{[DK]\to D_{s}^{*}\gamma}$, we need to calculate the scattering amplitude ^$\hat{\mathcal{M}}_{AB\to C\gamma}({\bf p},{\bf p}^{\prime})$. The corresponding Feynman diagrams are shown in Fig. 5.
The effective Lagrangians responsible for $\mathcal{D}^{(*)}\mathcal{D}^{(*)}V$ and $\mathcal{D}^{(*)}\mathcal{D}^{(*)}P$ interactions are Chen:2010re ; Chen:2014sra ; Faessler:2007gv
$\begin{split}\mathcal{L}_{\mathcal{D}^{(\ast)}\mathcal{D}^{(\ast)}\mathcal{P}}=&ig_{\mathcal{D}^{*}\mathcal{D}P}\mathcal{D}^{*\mu\dagger}\mathcal{D}\buildrel\leftrightarrow\over{\partial}_{\mu}\mathcal{P}\\ &-g_{\mathcal{D}^{*}\mathcal{D}^{*}P}\epsilon_{\mu\nu\alpha\beta}\mathcal{D}^{*\nu\dagger}(\partial^{\beta}\mathcal{D}^{*\alpha})\partial^{\mu}\mathcal{P}+H.c.,\\ \end{split}$ (52)
$\begin{split}\mathcal{L}_{\mathcal{D}^{(*)}\mathcal{D}^{(*)}\mathcal{V}}=&-ig_{\mathcal{DDV}}\mathcal{D}_{i}^{\dagger}\buildrel\leftrightarrow\over{\partial}_{\mu}\mathcal{D}^{j}(\mathcal{V}^{\mu})^{i}_{j}-2f_{\mathcal{D}^{*}\mathcal{D}\mathcal{V}}\epsilon_{\mu\nu\alpha\beta}\\ &\times(\partial^{\mu}\mathcal{V}^{\nu})^{i}_{j}(\mathcal{D}_{i}^{\dagger}\buildrel\leftrightarrow\over{\partial}{}^{\!\alpha}\mathcal{D}^{*\beta j}-\mathcal{D}_{i}^{*\beta\dagger}\buildrel\leftrightarrow\over{\partial}{}{\!{}^{\alpha}}\mathcal{D}^{j})\\ &+ig_{\mathcal{D}^{*}\mathcal{D}^{*}\mathcal{V}}\mathcal{D}^{*\nu\dagger}_{i}\buildrel\leftrightarrow\over{\partial}_{\!\mu}\mathcal{D}^{*j}_{\nu}(\mathcal{V}^{\mu})^{i}_{j}\\ &+4if_{\mathcal{D}^{*}\mathcal{D}^{*}\mathcal{V}}\mathcal{D}^{*\dagger}_{i\mu}(\partial^{\mu}\mathcal{V}^{\nu}-\partial^{\nu}\mathcal{V}^{\mu})^{i}_{j}\mathcal{D}^{*j}_{\nu},\end{split}$ (53)
[…]
$g_{D^{*}DP}$ $=$ $\frac{g}{f_{\pi}}\sqrt{m_{D^{*}}m_{D}},\quad g_{D^{*}D^{*}P}=\frac{2g}{f_{\pi}},\quad f_{D^{*}DV}=\frac{\lambda g_{V}}{\sqrt{2}},$
$f_{D^{*}D^{*}V}$ $=$ $\frac{\lambda g_{V}}{\sqrt{2}}m_{D^{*}},\quad g_{DDV}=g_{D^{*}D^{*}V}=\frac{\beta g_{V}}{\sqrt{2}},$ (54)
where the gauge couplings are: $\lambda=0.56$ Chen:2010re and $g=0.16$ Wang:2006ida .
The effective Lagrangians involving the electromagnetic interaction vertices are Faessler:2008vc ; Dong:2009uf
[…]
$\begin{split}\mathcal{L}_{\mathcal{D}^{*}\mathcal{D}\gamma}=&\bigg{\{}\frac{e}{4}g_{{D^{*+}D^{+}}\gamma}\varepsilon^{\mu\nu\alpha\beta}F_{\mu\nu}{D}^{*+}_{\alpha\beta}{D}^{-}\\ &+\frac{e}{4}g_{{D^{*0}{D}^{0}}\gamma}\varepsilon^{\mu\nu\alpha\beta}F_{\mu\nu}D^{*0}_{\alpha\beta}\bar{{D}}^{0}\bigg{\}}+H.c,\end{split}$ (56)
[…]
$\begin{split}\mathcal{L}_{\mathcal{K}^{*}\mathcal{K}\gamma}=&\bigg{\{}\frac{e}{4}g_{{K^{*+}K^{+}}\gamma}\varepsilon^{\mu\nu\alpha\beta}F_{\mu\nu}{K}^{*+}_{\alpha\beta}{K}^{-}\\ &+\frac{e}{4}g_{{K^{*0}{K}^{0}}\gamma}\varepsilon^{\mu\nu\alpha\beta}F_{\mu\nu}K^{*0}_{\alpha\beta}\bar{{K}}^{0}\bigg{\}}+H.c,\\ \end{split}$ (58)
[…]
with $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $D^{*0,+}_{\mu\nu}=\partial_{\mu}D^{*0,+}_{\nu}-\partial_{\nu}D^{*0,+}_{\mu}$. According to the experimental widths $\Gamma(D^{*+}\to D^{+}\gamma)=1.54$ keV ParticleDataGroup:2024cfk and $\Gamma(D^{*0}\to D^{0}\gamma)=26.04$ keV Dong:2008gb ; ParticleDataGroup:2024cfk , the coupling constants $g_{D^{*}D\gamma}$ are fixed as
$|g_{D^{*+}D^{+}\gamma}|$ $=$ $0.5\;\mathrm{GeV}^{-1},$ (60)
$|g_{D^{*0}D^{0}\gamma}|$ $=$ $2.0\;\mathrm{GeV}^{-1}.$ (61)
By using the experimental measurements of $\Gamma(K^{*+}\to K^{+}\gamma)=50.29\;\rm{keV}$ and $\Gamma(K^{*0}\to K^{0}\gamma)=116.19\;\rm{keV}$ ParticleDataGroup:2024cfk ; Faessler:2008vc , the $K^{*}K\gamma$ coupling constants are determined as
$|g_{K^{*+}K^{+}\gamma}|$ $=$ $0.836\;\mathrm{GeV}^{-1},$ (62)
$|g_{K^{*0}K^{0}\gamma}|$ $=$ $1.267\;\mathrm{GeV}^{-1}.$ (63)
In our calculation, we take $g_{D^{*+}D^{+}\gamma}=0.5$ GeV-1, $g_{D^{*0}D^{0}\gamma}=2.0$ GeV-1, $g_{K^{*+}K^{+}\gamma}=0.836$ GeV-1, and $g_{K^{*0}K^{0}\gamma}=-1.267$ GeV-1 as suggested in Ref. Faessler:2008vc . The interaction of four-point vertex $\mathcal{D}^{(*)}\mathcal{D}^{(*)}K\gamma$ can be obtained by replacing $\partial_{\mu}$ with $\partial_{\mu}+ieQA_{\mu}$ in Eq. (52).
Figure 5: Feynman diagrams for the radiative transition $D^{(*)}K\to D_{s}^{*}\gamma$.
With these effective Lagrangians, we can write down the scattering amplitudes for the $DK\to D_{s}^{*}\gamma$ process shown in Fig. 5,
$\begin{split}i\mathcal{M}_{a}^{A}=&i^{2}[ie\epsilon_{\gamma}^{*\mu}(-iq_{\mu}-ip_{1\mu)}][ig_{D^{*}DP}\epsilon_{D_{s}^{*}}^{*\nu}(-ip_{2\nu}+iq_{\nu})]\\ &\times\frac{i}{q^{2}-m_{D^{+}}^{2}},\end{split}$ (64)
$\begin{split}i\mathcal{M}_{b}^{A}=&i^{2}[ie\epsilon_{\gamma}^{*\mu}(-ip_{2\mu}-iq_{\mu})][ig_{D^{*}DP}\epsilon_{D_{s}^{*}}^{*\nu}(-iq_{\nu}+ip_{1\nu})]\\ &\times\frac{i}{q^{2}-M_{K^{+}}^{2}},\end{split}$ (65)
$\begin{split}i\mathcal{M}_{c}^{A}=&i^{2}\big{[}\frac{1}{4}eg_{D^{*0}D^{0}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}\big{[}-g_{D^{*}D^{*}P}\varepsilon_{\delta\phi\sigma\gamma}(-ip_{2}^{\delta})\\ &\times\epsilon_{D_{s}^{*}}^{*\phi}(-iq^{\gamma})\big{]}i\frac{-g^{\tau\sigma}+q^{\tau}q^{\sigma}/M_{D^{*0}}^{2}}{q^{2}-M_{D^{*0}}^{2}},\end{split}$ (66)
$\begin{split}i\mathcal{M}_{d}^{A}=&i^{2}\big{[}\frac{e}{4}eg_{D^{*+}D^{+}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}\big{[}-g_{D^{*}D^{*}P}\varepsilon_{\delta\phi\sigma\gamma}(-ip_{2}^{\delta})\\ &\times\epsilon_{D_{s}^{*}}^{*\phi}(-iq^{\gamma})\big{]}i\frac{-g^{\tau\sigma}+q^{\tau}q^{\sigma}/M_{D^{*+}}^{2}}{q^{2}-M_{D^{*+}}^{2}},\end{split}$ (67)
$\begin{split}i\mathcal{M}_{e}^{A}=&i^{2}\big{[}2f_{D^{*}DV}\varepsilon_{\delta\phi\sigma\gamma}(-iq^{\delta})(-ip_{4}^{\sigma}-ip_{1}^{\sigma})\epsilon_{D_{s}^{*}}^{*\gamma}\big{]}\\ &\times\big{[}\frac{1}{4}eg_{K^{*+}K^{+}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}i\frac{-g^{\phi\tau}+q^{\phi}q^{\tau}/M_{K^{*+}}^{2}}{q^{2}-M_{K^{*+}}^{2}},\end{split}$ (68)
$\begin{split}i\mathcal{M}_{f}^{A}=&i^{2}\big{[}2f_{D^{*}DV}\varepsilon_{\delta\phi\sigma\gamma}(-iq^{\delta})(-ip_{4}^{\sigma}-ip_{1}^{\sigma})\epsilon_{D_{s}^{*}}^{*\gamma}\big{]}\\ &\times\big{[}\frac{1}{4}eg_{K^{*0}K^{0}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}i\frac{-g^{\phi\tau}+q^{\phi}q^{\tau}/M_{K^{*0}}^{2}}{q^{2}-M_{K^{*0}}^{2}},\end{split}$ (69)
$\begin{split}i\mathcal{M}_{g}^{A}=&i^{2}[ig_{D^{*}DP}(-ip_{2\mu}+ip_{1\mu})](-ie)\\ &\times\big{[}g_{\alpha\beta}(-iq_{\nu}-ip_{4\nu})+g_{\nu\beta}(iq_{\alpha})+g_{\alpha\nu}(ip_{4\beta})\big{]}\\ &\times\epsilon_{\gamma}^{*\nu}\epsilon_{D_{s}^{*}}^{*\alpha}i\frac{-g^{\mu\beta}+q^{\mu}q^{\beta}/M_{D_{s}^{*+}}^{2}}{q^{2}-m_{D_{s}^{*+}}^{2}},\end{split}$ (70)
$\begin{split}i\mathcal{M}_{h}^{A}=&i^{2}[ig_{D^{*}DP}(-ip_{2\mu}+ip_{1\mu})](-ie)\\ &\times\big{[}g_{\alpha\beta}(-iq_{\nu}-ip_{4\nu})+g_{\nu\beta}(iq_{\alpha})+g_{\alpha\nu}(ip_{4\beta})\big{]}\\ &\times\epsilon_{\gamma}^{*\nu}\epsilon_{D_{s}^{*}}^{*\alpha}i\frac{-g^{\mu\beta}+q^{\mu}q^{\beta}/M_{D_{s}^{*+}}^{2}}{q^{2}-m_{D_{s}^{*+}}^{2}},\end{split}$ (71)
$i\mathcal{M}_{i}^{A}$ $=$ $i^{3}eg_{D^{*}DP}\epsilon_{\gamma}^{*\mu}\epsilon_{D_{s}^{*}\mu}^{*},$ (72)
$i\mathcal{M}_{l}^{A}$ $=$ $-i^{3}eg_{D^{*}DP}\epsilon_{\gamma}^{*\mu}\epsilon_{D_{s}^{*}\mu}^{*},$ (73)
where $\epsilon_{\gamma}$ and $\epsilon_{D_{s}^{*}}$ are the polarization vectors of photon and $D_{s}^{*}$ meson, respectively Chen:2024xlw , and $q$ is the momentum of the exchanged meson.
The total scattering amplitude for the $DK\to D_{s}^{*}\gamma$ process can be expressed as
$\begin{split}\mathcal{M}_{DK\to D_{s}^{*}\gamma}^{\rm Total}=\sum_{n=a}^{l}\mathcal{M}_{n}^{A}.\end{split}$ (74)
Considering the flavor wave function in Eq. (38), the decay amplitude in Eq. (49) need to be multiplied by the $1/\sqrt{2}$. To describe the off-shell effects of the exchanged mesons and account for the inner structures of the interaction vertices, we introduce the following form factor in the amplitudes Chen:2024xlw
$\mathcal{F}(\textbf{q})={\rm e}^{-\textbf{q}^{2n}/\Lambda^{\prime 2n}}.$ (75)
Meanwhile, we also calculate the decay process of $[D^{*}K]\to D_{s}^{*}\gamma$ similar to the $[DK]\to D_{s}^{*}\gamma$ one. Their corresponding Feynman amplitudes in Fig. 5 take the forms
$\begin{split}i\mathcal{M}_{a}^{B}=&i^{2}\big{[}\frac{1}{4}eg_{D^{*+}D^{+}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}\epsilon_{D^{*}}^{\tau}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(-ip_{1\alpha}g_{\beta\tau}+ip_{1\beta}g_{\alpha\tau})\big{]}\big{[}ig_{D^{*}DP}\epsilon_{D_{s}^{*}}^{*\phi}\\ &\times(-ip_{2\phi}+iq_{\phi})\big{]}\frac{i}{q^{2}-M_{D^{+}}^{2}},\end{split}$ (76)
$\begin{split}i\mathcal{M}_{b}^{B}=&i^{2}\big{[}ie\epsilon_{\gamma}^{*\mu}(-ip_{2\mu}-iq_{\mu})\big{]}\\ &\times\big{[}-g_{D^{*}D^{*}P}\epsilon_{\eta\tau\rho\sigma}(-iq^{\eta})\epsilon_{D_{s}^{*}}^{*\tau}\varepsilon_{D^{*}}^{\rho}(-ip_{1}^{\sigma})\big{]}\\ &\times\frac{i}{q^{2}-M_{K^{+}}^{2}},\end{split}$ (77)
$\begin{split}i\mathcal{M}_{c}^{B}=&i^{2}\big{[}\frac{1}{4}eg_{D^{*0}D^{0}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}\epsilon_{D^{*}}^{\tau}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(-ip_{1\alpha}g_{\beta\tau}+ip_{1\beta}g_{\alpha\tau})\big{]}\big{[}ig_{D^{*}DP}\epsilon_{D_{s}^{*}}^{*\phi}\\ &\times(-ip_{2\phi}+iq_{\phi})\big{]}\frac{i}{q^{2}-M_{D^{+}}^{2}},\end{split}$ (78)
$\begin{split}i\mathcal{M}_{d}^{B}=&i^{2}(-ie)\epsilon^{*\mu}_{\gamma}\epsilon^{\beta}_{D^{*}}\big{[}g_{\alpha\beta}(-iq_{\mu}-ip_{1\mu})+ig_{\mu\beta}(ip_{1\alpha})\\ &+g_{\mu\alpha}(iq_{\beta})\big{]}\big{[}-g_{D^{*}D^{*}P}\varepsilon_{\eta\tau\rho\sigma}(-ip_{2}^{\eta})\epsilon_{D_{s}^{*}}^{*\tau}(-iq^{\sigma})\big{]}\\ &\times i\frac{-g^{\alpha\rho}+q^{\alpha}q^{\rho}/{M_{D^{*+}}^{2}}}{q^{2}-M_{D^{*+}}^{2}},\end{split}$ (79)
$\begin{split}i\mathcal{M}_{e}^{B}=&i^{2}\big{[}ig_{D^{*}D^{*}V}\epsilon_{D_{s}^{*}}^{*\rho}\epsilon_{D^{*}}^{\phi}g_{\rho\phi}(-ip_{1\sigma}-ip_{4\sigma})\\ &+4if_{D^{*}D^{*}V}\epsilon_{D^{*}}^{\phi}\epsilon_{D_{s}^{*}}^{*\rho}(-iq_{\rho}g_{\phi\sigma}+iq_{\phi}g_{\rho\sigma})\big{]}\\ &\times\big{[}\frac{1}{4}eg_{K^{*}K\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}i\frac{-g^{\sigma\tau}+q^{\sigma}q^{\tau}/M_{K^{*+}}^{2}}{q^{2}-M_{K^{*+}}^{2}},\end{split}$ (80)
$\begin{split}i\mathcal{M}_{f}^{B}=&i^{2}\big{[}ig_{D^{*}D^{*}V}\epsilon_{D_{s}^{*}}^{*\rho}\epsilon_{D^{*}}^{\phi}g_{\rho\phi}(-ip_{1\sigma}-ip_{4\sigma})\\ &+4if_{D^{*}D^{*}V}\epsilon_{D^{*}}^{\phi}\epsilon_{D_{s}^{*}}^{*\rho}(-iq_{\rho}g_{\phi\sigma}+iq_{\phi}g_{\rho\sigma})\big{]}\\ &\times\big{[}\frac{1}{4}eg_{K^{*}K\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}i\frac{-g^{\sigma\tau}+q^{\sigma}q^{\tau}/M_{K^{*+}}^{2}}{q^{2}-M_{K^{*+}}^{2}},\end{split}$ (81)
$\begin{split}i\mathcal{M}_{g}^{B}=&i^{2}\big{[}-g_{D^{*}D^{*}P}\varepsilon_{\eta\tau\rho\sigma}(-ip_{2}^{\eta})\epsilon_{D^{*}}^{\rho}(-ip_{1}^{\sigma})\big{]}\\ &\times(-ie)[g_{\alpha\beta}(-ip_{4\mu}-iq_{\mu})+g_{\mu\beta}(iq_{\alpha})\\ &+g_{\mu\alpha}(ip_{4\beta})]\epsilon_{\gamma}^{*\mu}\epsilon_{D_{s}^{*}}^{*\alpha}i\frac{-g^{\tau\beta}+q^{\tau}q^{\beta}/M_{D_{s}^{*+}}^{2}}{q^{2}-M_{D_{s}^{*+}}^{2}},\end{split}$ (82)
$\begin{split}i\mathcal{M}_{h}^{B}=&i^{2}\big{[}-g_{D^{*}D^{*}P}\varepsilon_{\eta\tau\rho\sigma}(-ip_{2}^{\eta})\epsilon_{D^{*}}^{\rho}(-ip_{1}^{\sigma})\big{]}\\ &\times(-ie)[g_{\alpha\beta}(-ip_{4\mu}-iq_{\mu})+g_{\mu\beta}(iq_{\alpha})\\ &+g_{\mu\alpha}(ip_{4\beta})]\epsilon_{\gamma}^{*\mu}\epsilon_{D_{s}^{*}}^{*\alpha}i\frac{-g^{\tau\beta}+q^{\tau}q^{\beta}/M_{D_{s}^{*+}}^{2}}{q^{2}-M_{D_{s}^{*+}}^{2}},\end{split}$ (83)
$\begin{split}i\mathcal{M}_{i}^{B}=i^{2}eg_{D^{*}D^{*}P}\varepsilon_{\mu\nu\alpha\beta}(-ip_{2}^{\mu})\epsilon_{D_{s}^{*}}^{\nu}\epsilon_{D^{*}}^{\alpha}\epsilon_{\gamma}^{\beta},\end{split}$ (84)
$\begin{split}i\mathcal{M}_{l}^{B}=&i^{2}eg_{D^{*}D^{*}P}\varepsilon_{\mu\nu\alpha\beta}\epsilon_{\gamma}^{\mu}\epsilon_{D_{s}^{*}}^{\nu}\epsilon_{D^{*}}^{\alpha}(-ip_{1}^{\beta}).\end{split}$ (85)
In addition, we also calculate the decay process of $[D^{*}K]\to D_{s}\gamma$. The Feynman diagrams for the $D^{*}K\to D_{s}\gamma$ process are shown in Fig. 5, and their corresponding Feynman amplitudes are
$\begin{split}i\mathcal{M}_{a}^{C}=&i^{2}\big{[}ig_{D^{*}DP}(ip_{4\mu}+iq_{\mu})\epsilon_{D^{*}}^{\mu}\big{]}\big{[}ie\epsilon_{\gamma}^{*\nu}(-ip_{2\nu}-iq_{\nu})\big{]}\\ &\times\frac{i}{q^{2}-m_{K^{+}}^{2}},\end{split}$ (86)
$\begin{split}i\mathcal{M}_{b}^{C}=&i^{2}\big{[}ig_{D^{*}DP}(ip_{4\mu}+ip_{2\mu})]\big{]}(-ie)\epsilon^{*\rho}_{\gamma}\epsilon^{\beta}_{D^{*}}\\ \times&\big{[}g_{\alpha\beta}(-iq_{\rho}-ip_{1\rho})+ig_{\rho\beta}(ip_{1\alpha})+g_{\rho\alpha}(iq_{\beta})\big{]}\\ \times&i\frac{-g^{\mu\alpha}+q^{\mu}q^{\alpha}/{M_{D^{*+}}^{2}}}{q^{2}-M_{D^{*+}}^{2}},\end{split}$ (87)
$\begin{split}i\mathcal{M}_{c}^{C}=&i^{2}\big{[}-2f_{D^{*}DV}\varepsilon_{\delta\phi\sigma\gamma}(-iq^{\delta})(-ip_{4}^{\sigma}-ip_{1}^{\sigma})\epsilon_{D^{*}}^{\gamma}\big{]}\\ &\times\big{[}\frac{1}{4}eg_{K^{*+}K^{+}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}i\frac{-g^{\phi\tau}+q^{\phi}q^{\tau}/M_{K^{*+}}^{2}}{q^{2}-M_{K^{*+}}^{2}},\end{split}$ (88)
$\begin{split}i\mathcal{M}_{d}^{C}=&i^{2}\big{[}-2f_{D^{*}DV}\varepsilon_{\delta\phi\sigma\gamma}(-iq^{\delta})(-ip_{4}^{\sigma}-ip_{1}^{\sigma})\epsilon_{D^{*}}^{\gamma}\big{]}\\ &\times\big{[}\frac{1}{4}eg_{K^{*0}K^{0}\gamma}\varepsilon^{\mu\nu\alpha\beta}\epsilon_{\gamma}^{*\eta}(ip_{3\mu}g_{\nu\eta}-ip_{3\nu}g_{\mu\eta})\\ &\times(iq_{\alpha}g_{\beta\tau}-iq_{\beta}g_{\alpha\tau})\big{]}i\frac{-g^{\phi\tau}+q^{\phi}q^{\tau}/M_{K^{*0}}^{2}}{q^{2}-M_{K^{*0}}^{2}},\end{split}$ (89)
$\begin{split}i\mathcal{M}_{e}^{C}=&i^{2}\big{[}ig_{D^{*}DP}(ip_{2\mu}+iq_{\mu})\epsilon_{D^{*}}^{\mu}\big{]}\big{[}ie\epsilon_{\gamma}^{*\nu}(-ip_{4\nu}-iq_{\nu})\big{]}\\ &\times\frac{i}{q^{2}-m_{D_{s}^{+}}^{2}},\end{split}$ (90)
$\begin{split}i\mathcal{M}_{f}^{C}=&i^{2}\big{[}ig_{D^{*}DP}(ip_{2\mu}+iq_{\mu})\epsilon_{D^{*}}^{\mu}\big{]}\big{[}ie\epsilon_{\gamma}^{*\nu}(-ip_{4\nu}-iq_{\nu})\big{]}\\ &\times\frac{i}{q^{2}-m_{D_{s}^{+}}^{2}},\end{split}$ (91)
$\begin{split}i\mathcal{M}_{g}^{C}=-2ieg_{D^{*}DP}\epsilon_{\gamma}^{*\mu}\epsilon_{D^{*}\mu},\end{split}$ (92)
$\begin{split}i\mathcal{M}_{h}^{C}=-ieg_{D^{*}DP}\epsilon_{\gamma}^{*\mu}\epsilon_{D^{*}\mu}.\end{split}$ (93)
We have checked that gauge invariance is preserved in our amplitudes. That is, the total amplitude $\epsilon_{\gamma}^{*\mu}\mathcal{M}_{\mu}$ vanishes when we replace $\epsilon_{\gamma}^{*\mu}$ with the photon momentum $p_{3}^{\mu}$, i.e., $p_{3}^{\mu}\mathcal{M}_{\mu}=0$.
Based on heavy quark symmetry, our study can be easily extended to the $B^{(*)}\bar{K}$ system. Of course, the coupling constants should be replaced with the bottomed ones Oh:2000qr ; Xiao:2016mho ; Becirevic:2009yb ; Faessler:2008vc
$g_{B^{*}BP}=\frac{g}{f_{\pi}}\sqrt{m_{B^{*}}m_{B}},\quad g_{B^{*}B^{*}P}=\frac{2g}{f_{\pi}},\quad f_{B^{*}BV}=\frac{\lambda g_{V}}{\sqrt{2}},$
$f_{B^{*}B^{*}V}=\frac{\lambda g_{V}}{\sqrt{2}}m_{B^{*}},\quad g_{BBV}=g_{B^{*}B^{*}V}=\frac{\beta g_{V}}{\sqrt{2}},$
$g_{B^{*+}B^{+}\gamma}=0.5\;{\rm GeV}^{-1};\quad g_{B^{*0}B^{0}\gamma}=-1\;{\rm GeV}^{-1}.$ (94)
Figure 6: The dependence of the decay widths on the cutoff $\Lambda^{\prime}$ in the pure molecule picture.
Figure 7: The dependence of the decay widths on the cutoff $\Lambda^{\prime}$ with physical states as the mixtures of bare cores and molecules.
With the above preparations, we can calculate the radiative decay widths of the pure molecular states and the results are shown in Fig. 6. By taking the cutoff $\Lambda^{\prime}$ in the range 0.4--2.0 GeV, the decay width of $[DK]\to D^{*}_{s}\gamma$ is about 4--45 keV, and those of $[D^{*}K]\to D^{*}_{s}\gamma$ and $[D^{*}K]\to D_{s}\gamma$ are 0.4--2.1 keV and 1--16 keV, respectively. For the bottom sector, the decay widths of $[B\bar{K}]\to B_{s}^{*}\gamma$, $[B^{*}\bar{K}]\to B_{s}^{*}\gamma$, and $[B^{*}\bar{K}]\to B_{s}\gamma$ are about 8--238 keV, 0.2 keV--1.4 keV, and 4--250 keV, correspondingly. The ratio between $[DK]\to D_{s}^{*}\gamma$ and $[BK]\to B_{s}^{*}\gamma$ is $\frac{[BK]\to B_{s}^{*}\gamma}{[DK]\to D_{s}^{*}\gamma}\approx 2\sim 3$, which is close to the prediction of Ref.  Faessler:2008vc .
V.3 The radiative decay for the physical state
In the real situation, the physical state is a composite structure consisting of the bare quark-antiquark cores $N$ and hadron-hadron channels $[AB]$, and its radiative decay width can be then expressed as
$\begin{split}\Gamma_{N+[AB]\to C\gamma}^{\rm Total}=&\frac{1}{2J+1}\frac{\left|\bf p\right|}{32\pi^{2}m^{2}}\sum_{M}\int{\rm d}\Omega_{\textbf{p}}\\ &\times\left|\mathcal{M}_{{N}\to C\gamma}+\mathcal{M}_{{[AB]}\to C\gamma}\right|^{2}\\ =&\frac{1}{2J+1}\frac{\left|\bf p\right|}{32\pi^{2}m^{2}}\sum_{M}\int{\rm d}\Omega_{\textbf{p}}\bigg{(}\left|\mathcal{M}_{{N}\to C\gamma}\right|^{2}\\ &+\left|\mathcal{M}_{{[AB]}\to C\gamma}\right|^{2}+\mathcal{M}_{{N}\to C\gamma}^{\dagger}\mathcal{M}_{{[AB]}\to C\gamma}+H.c.\bigg{)},\\ \end{split}$ (95)
where $m$ is the mass of the physical state consisting of $N$ and $[AB]$. $\mathcal{M}_{{N}\to C\gamma}$ and $\mathcal{M}_{{[AB]}\to C\gamma}$ are the radiative transition amplitudes which already include the contributions from the probability amplitude or the wave function.
In Fig. 7, we show the radiative decay widths with the physical states as the mixings of bare cores and hadron-hadron channels. The ratio $\Gamma_{D_{s1}^{\prime}(2460)\to D_{s}\gamma}/\Gamma_{D_{s1}^{\prime}(2460)\to D_{s}^{*}\gamma}$ is bigger than 2 when $\Lambda^{\prime}$ is above 0.7 GeV, which is consistent with the current experimental data ParticleDataGroup:2024cfk . By comparing with Fig. 6 obtained by only including the pure molecular components, we notice that the widths change significantly when the bare cores are introduced, especially for $D_{s0}^{*}(2317)\to D_{s}^{*}\gamma$, $D_{s1}^{\prime}(2460)\to D_{s}\gamma$, $B_{s0}^{*}(5714)\to B_{s}^{*}\gamma$, $B_{s1}^{\prime}(5762)\to B_{s}^{*}\gamma$, and $B_{s1}^{\prime}(5762)\to B_{s}\gamma$.
Table 3: The radiative decay widths for two cases in units of keV. Here, case \@slowromancapi@ refers to results with the pure molecule assumption, and case \@slowromancapii@ is for the physical states as the mixings of bare cores and molecules.
Process $\Gamma[{D_{s0}^{*}\to D_{s}^{*}\gamma}]$ $\Gamma[D_{s1}^{\prime}\to D_{s}^{*}\gamma]$ $\Gamma[D_{s1}^{\prime}\to D_{s}\gamma]$ $\Gamma[B_{s0}^{*}\to B_{s}^{*}\gamma]$ $\Gamma[B_{s1}^{\prime}\to B_{s}^{*}\gamma]$ $\Gamma[B_{s1}^{\prime}\to B_{s}\gamma]$
3
5
7
9
11
13
$\Lambda^{\prime}$ (GeV) \@slowromancapi@ \@slowromancapii@ \@slowromancapi@ \@slowromancapii@ \@slowromancapi@ \@slowromancapii@ \@slowromancapi@ \@slowromancapii@ \@slowromancapi@ \@slowromancapii@ \@slowromancapi@ \@slowromancapii@
0.5 10.79 1.41 1.65 1.74 5.58 0.98 29.37 0.01 0.68 13.08 3.5 2.42
1.0 32.06 4.27 1.60 2.19 15.36 6.41 153.69 17.56 1.31 10.62 107.83 17.27
1.5 38.97 3.12 1.17 2.75 16.26 6.99 216.48 27.61 1.37 10.26 210.77 37.27
To highlight the role of the bare quark-antiquark cores, we discuss two cases here: case \@slowromancapi@ with the pure molecule assumption and case \@slowromancapii@ with the mixture picture. We summarize the numerical results for these two cases in Table 3. For $D_{s0}^{*}(2317)\to D_{s}^{*}\gamma$, the width in the case \@slowromancapi@ is about eight times more than that of case \@slowromancapii@. In the $D_{s1}^{\prime}(2460)\to D_{s}\gamma$ process, there is about a threefold change. In the $D_{s1}^{\prime}(2460)\to D_{s}^{*}\gamma$ process, the radiative decay width varies little with $\Lambda^{\prime}$ for both cases. In addition, the widths for the bottom sector change dozens of times with the introduction of bare $\bar{b}s$ cores, which indicates that the bore core components have a great influence.
Experimental data on these processes are still lacking, and different theoretical predictions with parity doubling model Bardeen:2003kt , effective Lagrangian approach Faessler:2007gv ; Faessler:2008vc ; Fu:2021wde , light-cone sum rules Colangelo:2005hv ; Wang:2008wz ; Wang:2006mf , and so on, vary in a very wide range. Ref. Wang:2019ehs treated $B_{s1}^{\prime}$ state as a molecular state and provided the widths of $B_{s1}^{\prime}\to B_{s}^{*}\gamma$ and $B_{s1}^{\prime}\to B_{s}\gamma$ as 0.4--2.6 and 45.2--79.8 keV, respectively, which agree relatively well with our results in case \@slowromancapi@.
The difference between different models is important for us to understand the inner structure of the $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ states and the predictions of their bottom analogs. Thus, we expect that the future experimental measurements on the radiative decays of $D_{s0}^{*}(2317)\to D_{s}^{*}\gamma$ and $D_{s1}^{\prime}(2460)\to D_{s}^{(*)}\gamma$, and their bottom-strange partners, are crucial for elucidating their inner structures.
VI Summary
In this work, we systematically study the $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom analogs within a sophisticated theoretical framework. We take the unquenched effect into account to investigate the low mass problem of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$. In our framework, the experiment masses of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ can be precisely reproduced, which indicates the unquenched effect plays an important role in the establishment of $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$.
In addition to the coupled channel effect between the bare $c\bar{s}$ core and $D^{(*)}K$ channel, we also consider the $D^{(*)}K$-$D^{(*)}K$ interaction. We notice that the further introduction of the latter interaction can cause a visible mass shift that can better accommodate the experimental measurements. Our results show that $D_{s0}^{*}(2317)$ and $D_{s1}^{\prime}(2460)$ are the mixtures of bare $c\bar{s}$ cores and $D^{(*)}K$ components, while the latter have more weights in the compositions. Furthermore, we predict the spectra of their bottom analogs using the same coupling constants and cutoff parameters. The predictions of their mass are $M_{B_{s0}^{*}}=5714$ MeV and $M_{B_{s1}^{\prime}}=5762$ MeV, respectively. They are the mixture of $\bar{b}s$ and $B^{(*)}\bar{K}$.
Due to the introduction of bare state core, the hadron-hadron channel binds tighter than the pure hadron-hadron system, which will affect their other properties. To decode the inner structure of physical states, we systematically investigate the radiative decay of $D_{s0}^{*}(2317)\to D_{s}^{*}\gamma$, $D_{s1}^{\prime}(2460)\to D_{s}^{*}\gamma$, $D_{s1}^{\prime}(2460)\to D_{s}\gamma$, $B_{s0}^{*}(5714)\to B_{s}^{*}\gamma$, $B_{s1}^{\prime}(5762)\to B_{s}^{*}\gamma$, and $B_{s1}^{\prime}(5762)\to B_{s}\gamma$ processes. Different components in the compositions give clearly different contributions to the radiative decay widths. Therefore we expect that future experimental measurements of these radiative decay processes will be definitely helpful to reveal their underlying structures.
Our exploration of the $D_{s0}^{*}(2317)$, $D_{s1}^{\prime}(2460)$, and their bottom analogs states provides a more profound understanding of the heavy-light meson system and nonperturbative QCD. We expect that our results can be further examined by future experiments.
ACKNOWLEDGMENTS
This project is partially supported by the National Natural Science Foundation of China under Grants No. 12175091, No. 12335001, No. 12247101, No. 12405098, No. 11975090, the 111 Project under Grant No. B20063, the innovation project for young science and technology talents of Lanzhou city under Grant No. 2023-QN-107, and the Science Foundation of Hebei Normal University with contract No. L2023B09.